### v1

# Rejection-Sampled Linear Codes for Lossy Compression and Channel Simulation

Abstract
We show that a linear code combined with rejection sampling can give a capacity-achieving scheme for simulating channels with additive noises with exchangeable distributions. Hence, it can be used in lossy source coding to achieve the rate-distortion function. Interestingly, unlike conventional linear covering codes for lossy compression which concerns the trade-off between the rate and the covering radius, our construction only requires the linear code to have a large distance (not a large covering radius), and is not sensitive to the rate of the linear code. Experiments reveal that our construction can outperform conventional covering codes for lossy source coding with Hamming distortion for a certain range of distortion levels, and performs well even when the blocklength is small (e.g., 24).
I Introduction
Channel simulation [1, 2, 3] concerns the setting where the encoder and decoder want to simulate a noisy channel via noiseless communication. One-shot channel simulation, where only one copy of the channel is simulated, has also been studied [4, 5, 6], where it was shown that the channel $p_{Y|X}$ with input distribution $p_{X}$ can be simulated using an expected $I(X;Y)+\log(I(X;Y)+1)+O(1)$ number of bits. While most previous channel simulation schemes utilize an unstructured approach [4, 5, 6, 7, 8, 9, 10], many of which have an exponential time complexity, it is also possible to construct schemes using structured codes which result in improved running time efficiency. For example, an additive noise channel over $\mathbb{R}^{n}$ can be simulated by combining rejection sampling and lattice codes [11]. Another example is the simulation of $n$ copies of the binary symmetric channels using the polar code, or using a more general linear code with a correction step [12]. A downside of the approach in [12] is that the blocklength $n$ must be large (the experiments in [12] are performed for $n\geq 2^{12}$). As a result, this approach can only be applied when the data is large, and requires a significant encoding and decoding delay if the encoding/decoding is performed in a streaming manner.
In this work, we propose combining greedy rejection sampling [4] and a linear code (which we call the inner code), yielding an efficient channel simulation scheme for small blocklengths. Our scheme can be applied to any additive noise channel with finite field vector input , i.e., $\mathbf{X}\in\mathbb{F}_{q}^{n}$, $\mathbf{Y}=\mathbf{X}+\mathbf{Z}$, where $\mathbf{Z}$ has an exchangeable distribution. We show that, as long as the inner code has a large enough distance, our scheme uses an amount of communication close to the capacity $C=\max_{p_{\mathbf{X}}}I(\mathbf{X};\mathbf{Y})$, and hence is capacity-achieving.
An application of our scheme is lossy source coding over $\mathbb{F}_{q}^{n}$ with Hamming distortion. By simulating the Hamming ball channel where $\mathbf{Z}$ is uniform over a Hamming ball, we can ensure that the distortion is always below the limit (as in $d$-semifaithful codes [13, 14, 15]), while attaining the rate-distortion function. Experiment results show that our scheme can outperform conventional linear covering codes [16, 17] in terms of rate-distortion performance over a certain range of distortion levels. Perhaps surprisingly, unlike conventional linear covering codes which requires a good trade-off between the rate and the covering radius, our scheme only requires the inner linear code to have a large distance (instead of a large covering radius). As long as the distance is large enough, our scheme can achieve the rate-distortion function, regardless of the rate of the inner code. This makes it significantly easier to find good inner codes for our scheme, giving a constructive approach to attain the rate-distortion function, compared to finding linear covering codes that attain the rate-distortion function where the only known techniques are non-constructive [16].
Another application is oblivious relaying [18, 19, 20], where the relay can apply channel simulation to compress the channel output. See Section II-C.
II Related Works
II-A Channel Simulation
It has been observed by Steiner [21] that rejection sampling [22] can be utilized to construct a one-shot channel simulation scheme. To maximize the acceptance probability at each iteration, Harsha et al. [4] proposed the greedy rejection sampling for discrete distributions. This approach was extended to general distributions by Flamich and Theis [9]. It was shown that the expected length of iteration number, i.e., the expected communication, is lower-bounded by $D_{\mathrm{KL}}(P\|Q)$ and upper bounded by $D_{\mathrm{KL}}(P\|Q)+2\log(D_{\mathrm{KL}}(P\|Q)+1)+O(1)$, where $P$ and $Q$ are the target and proposal distributions, respectively. By introducing a new quantity, the channel simulation divergence $D_{\mathrm{CS}}(\cdot\|\cdot)$, Goc and Flamich [23] further tightened the both bounds to $D_{\mathrm{CS}}(P\|Q)$ and $D_{\mathrm{CS}}(P\|Q)+\log(e+1)$, which implies the lower bound is achievable tightly. They also revealed that that for any causal rejection sampling scheme, the expected time complexity grows at least exponentially in the Rényi $\infty$-divergence $D_{\infty}(P\|Q)$ between $P$ and $Q$. To improve the running time, Flamich et al. [24, 10, 25] proposed an approach to partition the sample space so as to allow more efficient searching for the desired sample. However, this approach requires additional assumptions on $P$ and $Q$ to ensure efficiency. Note that the sampling schemes are unstructured. For practical implementation, it may be preferable to use structured approaches. Based on the capacity-achieving polar codes [26], Sriramu et al. [12] developed a simulation scheme for $n$ copies of symmetric binary-output channels, which shows asymptotic rate-optimality with time complexity $O(n\log n)$. In addition, they proposed using general linear codes with a correction step to simulate binary symmetric channels. Although both schemes approach the capacity for large $n$, their performance in short blocklength regimes is not guaranteed. Refer to [27] for an overview of various channel simulation schemes.
II-B Linear Codes for Lossy Source Coding
Using linear codes for lossy source coding with Hamming distortion has a long history. It was shown by Goblick [28] that linear codes can achieve the rate-distortion function. Due to the duality between channel coding and lossy source coding, many lossy source codes are designed based on good channel codes. The early attempt of Viterbi and Omura [29] showed that using trellis codes and the Viterbi algorithm [30] can approach the rate-distortion function as the trellis constraint length increases. However, the encoding complexity is exponential in the constraint length. After the low-density parity-check (LDPC) codes [31] attracted a lot of attention in channel coding, Matsunaga and Yamamoto [32] introduced them to lossy source coding. They showed that the ensemble of LDPC codes with variable degree $O(\log n)$ can attain the rate-distortion function asymptotically. Combined with the reinforced belief propagation [33], the LDPC coding scheme yields near optimal performance empirically with time complexity roughly $O(n\log n)$ [34, 35]. Another line of work considers using the dual code of LDPC codes, i.e., low-density generator matrix (LDGM) codes, for lossy source coding. Wainwright et al. [36] showed that the ensemble of check-regular LDGM codes approaches the rate-distortion function and the gap vanishes rapidly as the degree increases. Under their proposed message-passing algorithm with a decimation process [37], the LDGM codes yield performance close to the limit with practical complexity. The spatially coupled LDGM codes have also been applied to lossy source coding [38]. Furthermore, polar codes [26] were proven to achieve the rate-distortion function under the successive cancellation encoding with complexity $O(n\log n)$ [39].
II-C Oblivious Relaying
Another application is the information bottleneck channel, also known as the oblivious relay channel [18, 19, 20], which concerns the setting where the encoder encodes the message $M$ into an input sequence $\mathbf{X}=(X_{1},\ldots,X_{n})$, which is sent through a memoryless channel $p_{Y|X}$. A relay observes the output $\mathbf{Y}$ and compresses it into a description $W$ at a rate $R_{\mathrm{relay}}>0$. The decoder observes $W$ and attempts to decode $M$. The relay is required to be oblivious, that is, it cannot access the codebook used by the encoder and the decoder. The capacity is given by the information bottleneck [40] $\max_{p_{U|Y}:I(Y;U)\leq R_{\mathrm{relay}}}I(X;U)$ where $X\to Y\to U$ forms a Markov chain [18]. For example, consider $p_{Y|X}$ to be a binary symmetric channel $\mathrm{BSC}(\alpha)$. If the relay simulates the channel $p_{U|Y}=\mathrm{BSC}(\beta)$ so as to allow the decoder to recover an auxiliary output $\mathbf{U}$, the overall channel from $\mathbf{X}$ to $\mathbf{U}$ would be $\mathrm{BSC}(\alpha(1-\beta)+(1-\alpha)\beta)$. Hence, the encoder and decode can utilize standard coding schemes for the BSC on the auxiliary channel from $\mathbf{X}$ to $\mathbf{U}$ in order to achieve the capacity, and no specialized algorithm is needed to accommodate the oblivious relay. In comparison, if the relay utilized lossy source coding to compress $\mathbf{Y}$ (guaranteeing only a small Hamming distance between $\mathbf{Y}$ and $\mathbf{U}$) instead of channel simulation, then it would be difficult to compute the likelihood of $\mathbf{X}$ given $\mathbf{U}$. See Figure 1. Refer to [41] for the connection between oblivious relaying and channel simulation.
Figure 1: Top: An oblivious relay applying channel simulation to simulate a noisy channel $p_{U|Y}$. Bottom: The effect of channel simulation is that the simulated output $\mathbf{U}$ is distributed as if it is the output of $\mathbf{Y}$ passed through another noisy channel $p_{U|Y}$.
Notations
Logarithm is base $2$, and entropy is in bits. Write :=$\mathbb{N}:=\{1,2,\ldots\}$, $[a:b]:=\{a,\ldots,b\}$, $[n]:=\{1,\ldots,n\}$. For an event $E$, write $\mathbf{1}\{E\}$ for the indicator function which is $1$ if $E$ occurs, 00 otherwise. Write $\mathbb{F}_{q}$ for the finite field of order $q$. All vectors are assumed to be row vectors. For a vector $\mathbf{x}\in\mathbb{F}_{q}^{n}$, its Hamming weight is $\mathrm{wt}(\mathbf{x}):=|\{i\in[n]:\,x_{i}\neq 0\}|$.
Let $\mathcal{P}_{n}(\mathcal{X})$ be the set of probability mass functions over $\mathcal{X}$ where each entry is a multiple of $1/n$. For $\mathbf{x}=(x_{1},\ldots,x_{n})\in\mathcal{X}^{n}$, its type (or empirical distribution) is denoted as
$\mathrm{type}(\mathbf{x}):=(n^{-1}|\{i:\,x_{i}=y\}|)_{y\in\mathcal{X}}\in\mathcal{P}_{n}(\mathcal{X}).$
III The Channel Simulation Setting
In the setting of (one-shot) channel simulation with common randomness [4], there is a common randomness $B\sim p_{B}$ available to both the encoder and the decoder. The encoder observes an input $\mathbf{X}$, and sends a description $M=f(B,\mathbf{X})\in\mathcal{C}_{M}$ to the decoder, where $\mathcal{C}_{M}\subseteq\{0,1\}^{*}$ is a prefix-free codebook. The decoder then outputs $\hat{\mathbf{Y}}=g(B,M)$. The objective is to find a scheme $(p_{B},f,g)$ that ensures that $\hat{\mathbf{Y}}$ follows a prescribed conditional distribution $p_{\mathbf{Y}|\mathbf{X}}$ given $\mathbf{X}$ exactly or approximately, while minimizing the amount of communication $\mathbb{E}[|M|]$ (the expected length of $M$). In other words, we are simulating the noisy channel $p_{\mathbf{Y}|\mathbf{X}}$ using the noiseless communication of $M$. See Figure 2.
Figure 2: One-shot channel simulation.
Channel simulation is closely related to lossy source coding [42]. To compress $\mathbf{X}$ lossily under the distortion measure $d(\mathbf{x},\mathbf{y})$, we can perform channel simulation for a channel $p_{\mathbf{Y}|\mathbf{X}}$ which satisfies an expected distortion constraint $\mathbb{E}[d(\mathbf{X},\mathbf{Y})]\leq D$, or an almost-sure distortion constraint $\mathbb{P}(d(\mathbf{X},\mathbf{Y})\leq D)=1$ (e.g., in a $d$-semifaithful code [13]).
We are interested in simulating a finite field additive exchangeable noise (AXN) channel, where the channel input is a finite field vector $\mathbf{X}\in\mathbb{F}_{q}^{n}$, and the channel output is $\mathbf{Y}=\mathbf{X}+\mathbf{Z}$, where $\mathbf{Z}\in\mathbb{F}_{q}^{n}$ is a noise independent of $\mathbf{X}$. We require the entries of $\mathbf{Z}$ to form an exchangeable sequence, i.e., letting $p_{\mathbf{Z}}$ be the probability mass function of $\mathbf{Z}$, we have $p_{\mathbf{Z}}(\mathbf{z})=p_{\mathbf{Z}}(\mathbf{z}^{\prime})$ if $\mathbf{z}^{\prime}$ is formed by permuting the entries of $\mathbf{z}$, or equivalently, $p_{\mathbf{Z}}(\mathbf{z})$ depends only on $\mathrm{type}(\mathbf{z})$. In case if the field size is $q=2$, $\mathrm{type}(\mathbf{z})$ is equivalent to the Hamming weight $\mathrm{wt}(\mathbf{z})$, so $p_{\mathbf{Z}}(\mathbf{z})$ only depends on the weight. Note that memoryless additive noise channels, where the entries of $\mathbf{Z}$ are i.i.d., is a special case of AXN channels. We are particularly interested in the following AXN channels:
  • Memoryless $q$-ary symmetric channel, where $\mathbf{Z}=(Z_{i})_{i}$ with $Z_{i}$ being i.i.d. with $p_{Z_{i}}(0)=1-\alpha$, $p_{Z_{i}}(z)=\alpha/(q-1)$ for $z\neq 0$. Exactly simulating this channel ensures the expected distortion $\mathbb{E}[\mathrm{wt}(\mathbf{Y}-\mathbf{X})]\leq\alpha n$ for Hamming distortion $d(\mathbf{X},\mathbf{Y})=\mathrm{wt}(\mathbf{Y}-\mathbf{X})$.
  • Hamming ball channel, where $\mathbf{Z}\sim\mathrm{Unif}(\{\mathbf{z}:\,\mathrm{wt}(\mathbf{z})\leq w\})$ for a fixed $w\in[0:n]$. Exactly simulating this channel ensures the almost-sure distortion constraint $\mathrm{wt}(\mathbf{Y}-\mathbf{X})\leq w$.
  • Constant weight channel, where $\mathbf{Z}\sim\mathrm{Unif}(\{\mathbf{z}:\,\mathrm{wt}(\mathbf{z})=w\})$ for a fixed $w\in[0:n]$. Exactly simulating this channel also ensures the almost-sure distortion constraint.
IV Stochastic Syndrome Coder
Syndrome encoding and decoding are basic tools for linear codes. In this paper, we slightly generalizes the notion of syndrome coder by allowing a randomized output.
Definition 1.
An $(n,k)$-stochastic syndrome coder is characterized by a pair $(\mathbf{H},\kappa)$, where $\mathbf{H}\in\mathbb{F}_{q}^{(n-k)\times n}$ is a parity-check matrix with full row rank, and $\kappa$ is a conditional probability mass function from $\mathbb{F}_{q}^{n-k}$ to $\mathbb{F}_{q}^{n}$ satisfying that $\kappa(\mathbf{z}|\mathbf{s})=0$ (for $\mathbf{s}\in\mathbb{F}_{q}^{n-k}$, $\mathbf{z}\in\mathbb{F}_{q}^{n}$) if $\mathbf{z}\mathbf{H}^{\top}\neq\mathbf{s}$.
The concept of (stochastic) syndrome coder is general, and can be found in channel codes and lossy source codes:
  • Syndrome decoder for channel codes. The encoder encodes the message $\mathbf{M}\in\mathbb{F}_{q}^{k}$ into the codeword $\mathbf{X}=f(\mathbf{M})\in\mathbb{F}_{q}^{n}$ satisfying $\mathbf{X}\mathbf{H}^{\top}=\mathbf{0}$. The decoder observes $\mathbf{Y}\in\mathbb{F}_{q}^{n}$, generates $\hat{\mathbf{Z}}\in\mathbb{F}_{q}^{n}$ following the conditional distribution $\kappa(\hat{\mathbf{z}}|\mathbf{Y}\mathbf{H}^{\top})$, computes $\hat{\mathbf{X}}:=\mathbf{Y}-\hat{\mathbf{Z}}$, and finds the unique $\hat{\mathbf{M}}\in\mathbb{F}_{q}^{k}$ with $\hat{\mathbf{X}}=f(\hat{\mathbf{M}})$.
  • Syndrome encoder for lossy source codes. The encoder observes $\mathbf{X}\in\mathbb{F}_{q}^{n}$, generates $\hat{\mathbf{Z}}\in\mathbb{F}_{q}^{n}$ following $\kappa(\hat{\mathbf{z}}|-\mathbf{X}\mathbf{H}^{\top})$, computes $\hat{\mathbf{Y}}:=\mathbf{X}+\hat{\mathbf{Z}}$ (which must satisfy $\hat{\mathbf{Y}}\mathbf{H}^{\top}=\mathbf{0}$), and encodes $\hat{\mathbf{Y}}$ into $\mathbf{M}\in\mathbb{F}_{q}^{k}$. The decoder recovers $\hat{\mathbf{Y}}$ from $\mathbf{M}$ and outputs $\hat{\mathbf{Y}}$.
In both settings, the conditional distribution $\kappa$ is often taken to be the deterministic mapping which maps the syndrome $\mathbf{s}\in\mathbb{F}_{q}^{n-k}$ to the coset leader $\mathrm{argmin}_{\mathbf{z}\in\mathbb{F}_{q}^{n}:\,\mathbf{z}\mathbf{H}^{\top}=\mathbf{s}}\mathrm{wt}(\mathbf{z})$.
In this paper, we propose to use the stochastic syndrome coder for channel simulation. Compared to lossy source coding, we include a common randomness $(\boldsymbol{\Pi},\mathbf{b})$ to perform a random permutation and shift, so that the noise $\hat{\mathbf{z}}$ is uniform within any type class (a similar randomization is also present in [12]).
  • Stochastic syndrome encoder for channel simulation. Generate the common randomness $(\boldsymbol{\Pi},\mathbf{B})$ available to the encoder and the decoder, where $\boldsymbol{\Pi}\in\{0,1\}^{n\times n}$ is a uniformly random permutation matrix, and $\mathbf{B}\sim\mathrm{Unif}(\mathbb{F}_{q}^{n-k})$. [^1] The amount of common randmness needed to encode $(\boldsymbol{\Pi},\mathbf{b})$ is $\log(n!)+(n-k)\log q$ bits, which grows like $O(n\log n)$. The encoder observes $(\boldsymbol{\Pi},\mathbf{B})$ and $\mathbf{X}\in\mathbb{F}_{q}^{n}$, generates $\hat{\mathbf{V}}\in\mathbb{F}_{q}^{n}$ following the conditional distribution
    $\kappa(\hat{\mathbf{v}}\,|\,\mathbf{B}-\mathbf{X}\boldsymbol{\Pi}\mathbf{H}^{\top}),$
    computes
    $\hat{\mathbf{Z}}:=\hat{\mathbf{V}}\boldsymbol{\Pi}^{-1},\;\;\hat{\mathbf{Y}}:=\mathbf{X}+\hat{\mathbf{Z}},$
    which must satisfy $\hat{\mathbf{Y}}\boldsymbol{\Pi}\mathbf{H}^{\top}=\mathbf{B}$ (there are $q^{k}$ different $\hat{\mathbf{y}}$’s satisfying $\hat{\mathbf{y}}\boldsymbol{\Pi}\mathbf{H}^{\top}=\mathbf{B}$), and encodes $\hat{\mathbf{Y}}$ conditional on $(\boldsymbol{\Pi},\mathbf{B})$ into $\mathbf{M}\in\mathbb{F}_{q}^{k}$. The decoder recovers $\hat{\mathbf{Y}}$ from $\mathbf{M},\boldsymbol{\Pi},\mathbf{B}$ and outputs $\hat{\mathbf{Y}}$. For example, if $\mathbf{H}=[\tilde{\mathbf{H}}\,|\,\mathbf{I}_{n-k}]$ is in standard form, then the encoder can take $\mathbf{M}=(\hat{\mathbf{Y}}\boldsymbol{\Pi})_{1:k}$ to be the first $k$ entries of $\hat{\mathbf{Y}}\boldsymbol{\Pi}$, and the decoder can recover
    $\hat{\mathbf{Y}}=\big{[}\mathbf{M}\,\big{|}\,\mathbf{B}-\mathbf{M}\tilde{\mathbf{H}}^{\top}\big{]}\boldsymbol{\Pi}^{-1}.$ (1)
We now verify that the stochastic syndrome encoder simulates an additive exchangeable noise channel.
Proposition 2.
The stochastic syndrome encoder simulates an additive exchangeable noise channel, with noise distribution
$\mathbb{P}(\hat{\mathbf{Z}}=\hat{\mathbf{z}})=\mathbb{E}_{\boldsymbol{\Pi},\mathbf{S}}\left[\kappa(\hat{\mathbf{z}}\boldsymbol{\Pi}|\mathbf{S})\right],$
where $\boldsymbol{\Pi}\in\{0,1\}^{n\times n}$ is a uniformly random permutation matrix, independent of $\mathbf{S}\sim\mathrm{Unif}(\mathbb{F}_{q}^{n-k})$.
Proof:
We have
$\mathbb{P}(\hat{\mathbf{Z}}=\hat{\mathbf{z}}\,|\,\mathbf{X}=\mathbf{x})$ $=\mathbb{E}_{\boldsymbol{\Pi},\mathbf{B}}\left[\kappa(\hat{\mathbf{z}}\boldsymbol{\Pi}\,|\,\mathbf{B}-\mathbf{x}\boldsymbol{\Pi}\mathbf{H}^{\top})\right]$
$=\mathbb{E}_{\boldsymbol{\Pi},\mathbf{S}}\left[\kappa(\hat{\mathbf{z}}\boldsymbol{\Pi}|\mathbf{S})\right].$
Due to the random permutation, $\mathbb{P}(\hat{\mathbf{Z}}=\hat{\mathbf{z}})$ only depends on $\mathrm{type}(\hat{\mathbf{z}})$. ∎
V Rejection-Sampled Syndrome Encoder
A limitation of the stochastic syndrome encoder is that it is a fixed-length scheme, i.e., the length of the description $\mathbf{M}$ is always $k$. A fixed-length scheme can be suitable for approximate channel simulation where the actual noise simulated $\hat{\mathbf{Z}}$ only approximately follows the distribution of the desired noise $\mathbf{Z}$ (e.g., i.i.d. Bernoulli for binary symmetric channel), but is generally not suitable for exact channel simulation where $\hat{\mathbf{Z}}$ must have the same distribution as $\mathbf{Z}$ [27]. In this section, we will describe a class of variable-length schemes which combine the stochastic syndrome encoder with rejection sampling, suitable for exact simulation.
Definition 3.
A rejection-sampled syndrome encoder (RSSE) is characterized by a sequence of pairs $(\mathbf{H}_{i},\kappa_{i})_{i\in\mathbb{N}}$, where for $i\in\mathbb{N}$, $\mathbf{H}_{i}\in\mathbb{F}_{q}^{(n-k_{i})\times n}$ is a parity-check matrix with full row rank (we allow the $k_{i}$’s to be different), and $\kappa_{i}$ is a conditional probability mass function from $\mathbb{F}_{q}^{n-k_{i}}$ to $\mathbb{F}_{q}^{n}\cup\{\mathrm{e}\}$ satisfying that $\kappa_{i}(\mathbf{z}|\mathbf{s})=0$ (for $\mathbf{s}\in\mathbb{F}_{q}^{n-k_{i}}$, $\mathbf{z}\in\mathbb{F}_{q}^{n}$) if $\mathbf{z}\mathbf{H}_{i}^{\top}\neq\mathbf{s}$.
We call the sequence of codes $(\mathbf{H}_{i})_{i}$ the inner code sequence. Definition 3 describes a general class of schemes instead of one specific scheme. The construction of a specific capacity-achieving RSSE will be discussed in Section VI. Note that each $(\mathbf{H}_{i},\kappa_{i})$ in an RSSE is almost a stochastic syndrome encoder, except that $\kappa_{i}$ is allowed to output a “rejection symbol” $\mathrm{e}$ which signals that we should reject the current iteration and proceed to $(\mathbf{H}_{i+1},\kappa_{i+1})$. To perform encoding in the RSSE, the encoder performs the same steps as in the stochastic syndrome encoder in Section IV on $(\mathbf{H}_{i},\kappa_{i})$ for $i=1,2,\ldots$ (each with an independent common randomness $(\boldsymbol{\Pi}_{i},\mathbf{B}_{i})$) until the first time where $\hat{\mathbf{V}}_{i}\neq\mathrm{e}$ (let it be iteration $L\in\mathbb{N}$), and selects $\hat{\mathbf{Z}}=\hat{\mathbf{V}}_{L}\boldsymbol{\Pi}_{L}^{-1}$, $\hat{\mathbf{Y}}=\mathbf{X}+\hat{\mathbf{Z}}$. The encoder then encodes $\hat{\mathbf{Y}}$ conditional on $(\boldsymbol{\Pi}_{L},\mathbf{B}_{L})$ into $\mathbf{M}\in\mathbb{F}_{q}^{k_{L}}$, and sends $(L,\mathbf{M})$. Practically, the encoder encodes $L$ using the Huffman code [43] or Elias codes [44], and then append $\mathbf{M}\in\mathbb{F}_{q}^{k_{L}}$ to form the transmitted bit sequence, which is prefix-free since the decoder can recover $L$ and then know the length of $\mathbf{M}$ which is $k_{L}$. The decoder can then decode $\mathbf{M}$ into $\hat{\mathbf{Y}}$ conditional on $(\boldsymbol{\Pi}_{L},\mathbf{B}_{L})$. If the Huffman code is used, the expected communication cost is at most $H(L)+\mathbb{E}[k_{L}\log q]+1$.
The precise algorithm for RSSE is given in Algorithm 1, where we assume there are two synchronized random number generators (RNG) available to the encoder and the decoder which can be used to generate the common randomness $(\boldsymbol{\Pi}_{i},\mathbf{B}_{i})$. In practice, they will initialize two pseudo-random number generators with the same seed.
Procedure $\textsc{Encode}((\mathbf{H}_{i},\kappa_{i})_{i\in\mathbb{N}},\mathbf{X},\mathfrak{G}):$
    Input: rejection-sampled syndrome encoder $(\mathbf{H}_{i},\kappa_{i})_{i\in\mathbb{N}}$,
            input $\mathbf{X}\in\mathbb{F}_{q}^{n}$, RNG $\mathfrak{G}$
    Output: iteration $L\in\mathbb{N}$, description $\mathbf{M}\in\mathbb{F}_{q}^{k_{L}}$
1:for $i=1,2,\ldots$ do
2:     Generate permutation matrix $\boldsymbol{\Pi}_{i}\in\{0,1\}^{n\times n}$
3:           and $\mathbf{B}_{i}\sim\mathrm{Unif}(\mathbb{F}_{q}^{n-k_{i}})$ using $\mathfrak{G}$
4:     Generate $\hat{\mathbf{V}}_{i}\in\mathbb{F}_{q}^{n}\cup\{\mathrm{e}\}$ using local randomness
5:           following distribution $\kappa_{i}(\hat{\mathbf{v}}\,|\,\mathbf{B}_{i}-\mathbf{X}\boldsymbol{\Pi}_{i}\mathbf{H}_{i}^{\top})$
6:     if $\hat{\mathbf{V}}_{i}\neq\mathrm{e}$ then
7:         $\hat{\mathbf{Z}}\leftarrow\hat{\mathbf{V}}_{i}\boldsymbol{\Pi}_{i}^{-1}$,  $\hat{\mathbf{Y}}\leftarrow\mathbf{X}+\hat{\mathbf{Z}}$
8:         Encode $\hat{\mathbf{Y}}$ conditional on $(\boldsymbol{\Pi}_{i},\mathbf{B}_{i})$ into $\mathbf{M}\in\mathbb{F}_{q}^{k_{L}}$
9:               (e.g., using (1) if $\mathbf{H}_{i}$ is in standard form)
10:         return $(i,\mathbf{M})$
11:     end if
12:end for
Procedure $\textsc{Decode}((\mathbf{H}_{i},\kappa_{i})_{i\in\mathbb{N}},L,\mathbf{M},\mathfrak{G}):$
    Input: $(\mathbf{H}_{i},\kappa_{i})_{i\in\mathbb{N}}$, $L\in\mathbb{N}$, $\mathbf{M}\in\mathbb{F}_{q}^{k_{L}}$, RNG $\mathfrak{G}$
    Output: output $\mathbf{X}\in\mathbb{F}_{q}^{n}$
1:for $i=1,\ldots,L$ do
2:     Generate permutation matrix $\boldsymbol{\Pi}_{i}\in\{0,1\}^{n\times n}$
3:           and $\mathbf{B}_{i}\sim\mathrm{Unif}(\mathbb{F}_{q}^{n-k_{i}})$ using $\mathfrak{G}$
4:end for
5:Decode $\hat{\mathbf{Y}}$ conditional on $(\boldsymbol{\Pi}_{L},\mathbf{B}_{L})$ from $\mathbf{M}$
6:return $\hat{\mathbf{Y}}$
Algorithm 1 Rejection-sampled syndrome encoder
We can check that RSSE simulates an additive exchangeable noise channel.
Proposition 4.
The rejection-sampled syndrome encoder simulates an additive exchangeable noise channel, with noise distribution
$\mathbb{P}(\hat{\mathbf{Z}}=\hat{\mathbf{z}})=\sum_{i=1}^{\infty}\mathbb{P}(L\geq i)\mathbb{E}\left[\kappa_{i}(\hat{\mathbf{z}}\boldsymbol{\Pi}_{i}\,|\,\mathbf{S}_{i})\right]$
where $\boldsymbol{\Pi}_{i}\in\{0,1\}^{n\times n}$ is a uniformly random permutation matrix, independent of $\mathbf{S}_{i}\sim\mathrm{Unif}(\mathbb{F}_{q}^{n-k_{i}})$, where $L\in\mathbb{N}$ is the index selected by RSSE, with $\mathbb{P}(L\geq i)$ given recursively as
$\mathbb{P}(L\geq i+1)=\mathbb{P}(L\geq i)\mathbb{E}\left[\kappa_{i}(\mathrm{e}\,|\,\mathbf{S}_{i})\right].$
Proof:
At iteration $i$ (i.e., conditional on $L\geq i$), using the same arguments as Proposition 2, the encoder selects $\hat{\mathbf{z}}$ with probability
$\mathbb{P}(\hat{\mathbf{Z}}=\hat{\mathbf{z}},\,L=i\,|\,L\geq i)=\mathbb{E}\left[\kappa_{i}(\hat{\mathbf{z}}\boldsymbol{\Pi}_{i}\,|\,\mathbf{S}_{i})\right].$
Also, the probability of rejection at iteration $i$ is
$\mathbb{P}(L\geq i+1\,|\,L\geq i)=\mathbb{E}\left[\kappa_{i}(\mathrm{e}\,|\,\mathbf{S}_{i})\right].$
VI Greedy Rejection-Sampled Syndrome Encoder
In this section, we will give an algorithm for the construction of an RSSE for the exact simulation of an additive exchangeable noise channel with noise distribution $p_{\mathbf{Z}}(\mathbf{z})$, called greedy rejection-sampled syndrome encoder (GRSSE). It is inspired by the greedy rejection sampling scheme [4], which maximizes the probability of acceptance at iteration $1$, then maximizes the probability of acceptance at iteration $2$, and so on. Unlike greedy rejection sampling, we do not have a fixed proposal distribution, but rather has the flexibility to design $\kappa_{i}$ which helps us further maximize the probability of acceptance.
Definition 5.
Given a sequence of parity-check matrices $(\mathbf{H}_{i})_{i}$ and the target exchangeable noise distribution $p_{\mathbf{Z}}$, the greedy rejection-sampled syndrome encoder (GRSSE) $(\mathbf{H}_{i},\kappa_{i})_{i}$ is defined recursively as $p_{\mathbf{Z}}^{(1)}(\mathbf{z})=p_{\mathbf{Z}}(\mathbf{z})$ for $\mathbf{z}\in\mathbb{F}_{q}^{n}$, and for $i=1,2,\ldots$, using the notations in Proposition 4,
$\kappa_{i}=\underset{\kappa^{\prime}_{i}:\,\mathbb{E}[\kappa^{\prime}_{i}(\mathbf{z}\boldsymbol{\Pi}_{i}\,|\,\mathbf{S}_{i})]\leq p_{\mathbf{Z}}^{(i)}(\mathbf{z}),\,\forall\mathbf{z}\in\mathbb{F}_{q}^{n}}{\mathrm{argmin}}\mathbb{E}[\kappa^{\prime}_{i}(\mathrm{e}|\mathbf{S}_{i})],$ (2)
where the argmin is over $\kappa^{\prime}_{i}$ satisfying Definition 3 (if the argmin is not unique, select $\kappa_{i}$ arbitrarily), and
$p_{\mathbf{Z}}^{(i+1)}(\mathbf{z})=\frac{p_{\mathbf{Z}}^{(i)}(\mathbf{z})-\mathbb{E}[\kappa_{i}(\mathbf{z}\boldsymbol{\Pi}_{i}\,|\,\mathbf{S}_{i})]}{\mathbb{E}[\kappa_{i}(\mathrm{e}\,|\,\mathbf{S}_{i})]}.$
We can terminate this process when $\mathbb{E}[\kappa_{i}(\mathrm{e}\,|\,\mathbf{S}_{i})]=0$.
Intuitively, GRSSE first selects $\kappa_{1}$ that minimizes $\mathbb{E}[\kappa_{1}(\mathrm{e}\,|\,\mathbf{S}_{1})]$, the probability of rejection at iteration $1$ (see Proposition 4). Note that we require $\mathbb{E}[\kappa_{1}(\mathbf{z}\boldsymbol{\Pi}_{1}\,|\,\mathbf{S}_{1})]=\mathbb{P}(\hat{\mathbf{Z}}=\mathbf{z},\,L=1)\leq p_{\mathbf{Z}}(\mathbf{z})$ if we want to exactly simulate $p_{\mathbf{Z}}$, or else the probability of the noise $\mathbf{z}$ will be too large. Hence, we minimize $\mathbb{E}[\kappa_{1}(\mathrm{e}\,|\,\mathbf{S}_{1})]$ subject to $\mathbb{E}[\kappa_{1}(\mathbf{z}\boldsymbol{\Pi}_{1}\,|\,\mathbf{S}_{1})]\leq p_{\mathbf{Z}}(\mathbf{z})$ for all $\mathbf{z}$. The gaps $p_{\mathbf{Z}}(\mathbf{z})-\mathbb{E}[\kappa_{1}(\mathbf{z}\boldsymbol{\Pi}_{1}\,|\,\mathbf{S}_{1})]$, normalized to be a probability distribution $p_{\mathbf{Z}}^{(2)}(\mathbf{z})=p_{\mathbf{Z}|L\geq 2}(\mathbf{z})$, must be filled in the remaining iterations to ensure exact simulation. We then continue this process for iteration $2$ to simulate the noise distribution $p_{\mathbf{Z}}^{(2)}(\mathbf{z})$, and so on. It can be checked that $p_{\mathbf{Z}}^{(i)}(\mathbf{z})=p_{\mathbf{Z}|L\geq i}(\mathbf{z})$.
We now describe the computation of the argmin in (2). There are various symmetries in (2) that allow us to reduce the dimension of the problem. First, due to the random permutation $\boldsymbol{\Pi}_{i}$, the distributions $p_{\mathbf{Z}}^{(i)}(\mathbf{z})$ are always exchangeable, and hence we only need to keep track of the distribution of the type $\mathbf{p}=\mathrm{type}(\mathbf{z})$, i.e., we consider $p_{\mathbf{P}}^{(i)}(\mathbf{p})=\sum_{\mathbf{z}:\mathrm{type}(\mathbf{z})=\mathbf{p}}p_{\mathbf{Z}}^{(i)}(\mathbf{z})$ for the distribution of $\mathbf{P}=\mathrm{type}(\mathbf{Z})\in\mathcal{P}_{n}(\mathbb{F}_{q})$ when $\mathbf{Z}\sim p_{\mathbf{Z}}^{(i)}$, and consider a conditional distribution $\tilde{\kappa}_{i}(\mathbf{p}|\mathbf{s})$ from the syndrome to the type instead of $\kappa_{i}(\mathbf{z}|\mathbf{s})$. This significantly reduces the dimension of the problem since $|\mathcal{P}_{n}(\mathbb{F}_{q})|=O(n^{q-1})\ll|\mathbb{F}_{q}^{n}|=q^{n}$. Second, the constraint in Definition 3 requires $\tilde{\kappa}_{i}(\mathbf{p}|\mathbf{s})>0$ only if ther exists $\mathbf{z}$ with $\mathrm{type}(\mathbf{z})=\mathbf{p}$ and $\mathbf{z}\mathbf{H}_{i}^{\top}=\mathbf{s}$. Define the type set of the syndrome $\mathbf{s}$ to be
$t_{\mathbf{H}}(\mathbf{s})=\big{\{}\mathrm{type}(\mathbf{z}):\,\mathbf{z}\in\mathbb{F}_{q}^{n},\,\mathbf{z}\mathbf{H}^{\top}=\mathbf{s}\big{\}}\in 2^{\mathcal{P}_{n}(\mathbb{F}_{q})}.$
Then syndromes with the same type set are subject to the same constraint, and hence can be considered together. The number of different type sets is often significantly smaller than the number of syndromes. For example, the $[24,12,8]$ Golay code [45] has $2^{12}$ syndromes but only $5$ type sets. Define the type set distribution to be the probability mass function
$p_{t_{\mathbf{H}}}(t):=q^{-(n-k)}|\{\mathbf{s}:\,t_{\mathbf{H}}(\mathbf{s})=t\}|$ (3)
over $t\in t_{\mathbf{H}}(\mathbb{F}_{q}^{n-k})$, i.e., it is the distribution of $T=t_{\mathbf{H}}(\mathbf{S})$ where $\mathbf{S}\sim\mathrm{Unif}(\mathbb{F}_{q}^{n-k})$. For $t\in t_{\mathbf{H}_{i}}(\mathbb{F}_{q}^{n-k})$, $\mathbf{p}\in\mathcal{P}_{n}(\mathbb{F}_{q})$, let
$\gamma_{i,t,\mathbf{p}}:=q^{-(n-k)}\sum_{\mathbf{s},\mathbf{z}:t_{\mathbf{H}}(\mathbf{s})=t,\,\mathrm{type}(\mathbf{z})=\mathbf{p}}\kappa_{i}(\mathbf{z}|\mathbf{s}).$
Note that $\sum_{t,\mathbf{p}}\gamma_{i,t,\mathbf{p}}$ is generally not $1$ since the random mapping $\kappa_{i}$ may have a positive probability of mapping $\mathbf{s}$ to the erasure symbol $\mathrm{e}$ which indicates a rejection. We can maximize the probability of acceptance by maximizing $\sum_{t,\mathbf{p}}\gamma_{i,t,\mathbf{p}}$. Hence, the problem (2) is equivalent to the problem of maximizing $\sum_{t,\mathbf{p}}\gamma_{i,t,\mathbf{p}}$ subject to the linear constraints
$\begin{array}[]{ll}\gamma_{i,t,\mathbf{p}}\geq 0,\;\forall t,\mathbf{p};&\sum_{\mathbf{p}}\gamma_{i,t,\mathbf{p}}\leq p_{t_{\mathbf{H}_{i}}}(t),\;\forall t;\\ \sum_{t}\gamma_{i,t,\mathbf{p}}\leq p_{\mathbf{P}}^{(i)}(\mathbf{p}),\;\forall\mathbf{p};&\gamma_{i,t,\mathbf{p}}=0\;\;\text{if}\;\mathbf{p}\notin t.\end{array}$
This maximization can be solved efficiently using a linear programming or maximum flow algorithm. Refer to Algorithm 2, which is for the precomputation of $\kappa_{i}$ in GRSSE. In practice, before the communication commences, the encoder first precomputes the $\kappa_{i}$’s using Algorithm 2. During the communication, the encoder and decoder use Algorithm 1 to perform encoding and decoding. Although Algorithm 2 may take a significant amount of time, it is only computed once, and the $\kappa_{i}$’s can be reused for subsequent communications. Algorithm 2 also computes the distribution $p_{L}$, which is useful if we are going to encode $L$ via Huffman coding.
[…]
    Input: type distribution $p_{\mathbf{P}}(\mathbf{p})=\sum_{\mathbf{z}:\mathrm{type}(\mathbf{z})=\mathbf{p}}p_{\mathbf{Z}}(\mathbf{z})$,
            parity-check matrices $(\mathbf{H}_{i})_{i}$
    Output: RSSE $(\mathbf{H}_{i},\kappa_{i})_{i}$, distribution $p_{L}$
1:$p_{\mathbf{P}}^{(1)}(\mathbf{p})\leftarrow p_{\mathbf{P}}(\mathbf{p})$
2:$S\leftarrow 1$
3:for $i=1,2,\ldots$ do
4:     $p_{t_{\mathbf{H}_{i}}}(t)\leftarrow q^{-(n-k)}|\{\mathbf{s}:\,t_{\mathbf{H}}(\mathbf{s})=t\}|$, $\forall t\in t_{\mathbf{H}_{i}}(\mathbb{F}_{q}^{n-k})$
5:     Construct weighted directed graph $G$ with
6:        vertices $V=\{\mathrm{src},\mathrm{des}\}\cup t_{\mathbf{H}_{i}}(\mathbb{F}_{q}^{n-k})\cup\mathcal{P}_{n}(\mathbb{F}_{q})$,
7:        edges $(t,\mathbf{p})$ with capacity $\infty$ for $t,\mathbf{p}$ with $\mathbf{p}\in t$,
8:        edges $(\mathrm{src},t)$ with cap. $p_{t_{\mathbf{H}_{i}}}(t)$ for $t\in t_{\mathbf{H}_{i}}(\mathbb{F}_{q}^{n-k})$,
9:        edges $(\mathbf{p},\mathrm{des})$ with cap. $p_{\mathbf{P}}^{(i)}(\mathbf{p})$ for $\mathbf{p}\in\mathcal{P}_{n}(\mathbb{F}_{q})$
10:     Compute maximum flow of $G$ from $\mathrm{src}$ to $\mathrm{des}$
11:     $F\leftarrow$ total flow, $\;\gamma_{i,t,\mathbf{p}}\leftarrow$ flow along edge $(t,\mathbf{p})$
12:     $\kappa_{i}(\mathbf{z}|\mathbf{s})\leftarrow 0$, $\forall\mathbf{s},\mathbf{z}$
13:     $\mathrm{vis}(\mathbf{s},\mathbf{p})\leftarrow 0$, $\forall\mathbf{s},\mathbf{p}$
14:     for $\mathbf{z}\in\mathbb{F}_{q}^{n}$ do
15:         $\mathbf{s}\leftarrow\mathbf{z}\mathbf{H}_{i}^{\top}$, $\;t\leftarrow t_{\mathbf{H}_{i}}(\mathbf{s}),$ $\;\mathbf{p}\leftarrow\mathrm{type}(\mathbf{z})$
16:         if $\mathrm{vis}(\mathbf{s},\mathbf{p})=0$ then
17:              $\kappa_{i}(\mathbf{z}|\mathbf{s})\leftarrow\gamma_{i,t,\mathbf{p}}/p_{t_{\mathbf{H}_{i}}}(t)$, $\;\mathrm{vis}(\mathbf{s},\mathbf{p})\leftarrow 1$
18:         end if
19:     end for
20:     $\kappa_{i}(\mathrm{e}|\mathbf{s})\leftarrow 1-\sum_{\mathbf{p}\in t}\gamma_{i,t,\mathbf{p}}/p_{t_{\mathbf{H}_{i}}}(t)$ (where $t=t_{\mathbf{H}_{i}}(\mathbf{s})$)
21:     $p_{L}(i)\leftarrow SF$
22:     if $F=1$ then
23:         return $(\mathbf{H}_{j},\kappa_{j})_{j\in[i]}$, $p_{L}$
24:     end if
25:     $p_{\mathbf{P}}^{(i+1)}(\mathbf{p})\leftarrow(p_{\mathbf{P}}^{(i)}(\mathbf{p})-\sum_{\mathbf{s}}\gamma_{i,\mathbf{s},\mathbf{p}})/(1-F)$
26:     $S\leftarrow S(1-F)$
27:end for
Algorithm 2 Precomputation for greedy rejection-sampled syndrome encoder
VII Theoretical Analysis of GRSSE
We now analyze the performance of GRSSE. For simplicity, we consider the inner codes $\mathbf{H}_{i}=\mathbf{H}$ to be the same for all $i$. We call this the $\mathbf{H}$-GRSSE or pure GRSSE. Let the distance of the code be $d:=\min_{\mathbf{z}\neq 0:\,\mathbf{z}\mathbf{H}^{\top}=0}\mathrm{wt}(\mathbf{z})$. Theorem 6 shows that $\mathbf{H}$-GRSSE achieves the capacity under very mild conditions. Note that the bound in Theorem 6 holds regardless of the distribution of the input $\mathbf{X}$. A proof sketch will be presented later in this section. The complete proof is in Appendix -B.
Theorem 6.
The $\mathbf{H}$-GRSSE for the simulation of the additive exchangeable noise channel with noise distribution $p_{\mathbf{Z}}$, with $\mathbf{H}_{i}=\mathbf{H}$ where the code $\mathbf{H}\in\mathbb{F}_{q}^{(n-k)\times n}$ has a distance $d$, gives a number of iteration $L$ satisfying
$\mathbb{E}[\log L]$ $\leq C-\mathbb{P}\left(\mathrm{wt}(\mathbf{Z})<d/2\right)k\log q+\eta+1,$ (4)
where $C=I(\mathbf{X};\mathbf{Y})$ is the channel capacity (where $\mathbf{X}\sim\mathrm{Unif}(\mathbb{F}_{q}^{n})$, $\mathbf{Z}\sim p_{\mathbf{Z}}$, $\mathbf{Y}=\mathbf{X}+\mathbf{Z}$) and $\eta:=(1+e^{-1})\log e$. Hence, the expected number of bits of communication (for sending $L,\mathbf{M}$) is at most
$C+\mathbb{P}\left(\mathrm{wt}(\mathbf{Z})\geq d/2\right)k\log q$
$\;+\log\left(C-\mathbb{P}\left(\mathrm{wt}(\mathbf{Z})<d/2\right)k\log q+\eta+2\right)+\eta+3.$ (5)
In particular, if $\mathrm{wt}(\mathbf{Z})<d/2$ almost surely (e.g., the Hamming ball channel and constant weight channel in Section III), the communication is upper-bounded by
$C+\log\left(C-k\log q+\eta+2\right)+\eta+3.$
To achieve the capacity $C$, we need $d$ to be large enough so $\mathrm{wt}(\mathbf{Z})<d/2$ with high probability. A direct corollary is that for every sequence of AXN channels with noise distributions $p_{\mathbf{Z}^{(n)}}$ over $\mathbb{F}_{q}^{n}$ and capacities $C^{(n)}$ for $n=1,2,\ldots$, with asymptotic capacity $\overline{C}=\lim_{n\to\infty}C^{(n)}/n$, as long as we find a sequence of inner codes $\mathbf{H}^{(n)}$ with distance $d^{(n)}$ such that $\mathbb{P}(\mathrm{wt}(\mathbf{Z}^{(n)})\geq d^{(n)}/2)\to 0$, then GRSSE achieves the asymptotic capacity, in the sense that the expected number of bits of communication per dimension $n$ approaches $\overline{C}$. As a result, by simulating the sequence of Hamming ball channels $p_{\mathbf{Z}^{(n)}}=\mathrm{Unif}(\{\mathbf{z}\in\mathbb{F}_{q}^{n}:\,\mathrm{wt}(\mathbf{z})\leq\alpha n\})$ with $\overline{C}=\log q+(1-\alpha)\log(1-\alpha)+\alpha\log(\alpha/(q-1))$ for $\alpha\leq 1-1/q$, as long as the distances satisfy $\mathrm{liminf}_{n}d^{(n)}/n>2\alpha$, GRSSE can achieve the rate-distortion function for Hamming distortion.
Interestingly, unlike conventional lossy source coding which requires a small covering radius of the code, GRSSE requires a large enough distance $d$. GRSSE does not require $k$ and $d$ to be precise or optimal in any sense, as $d$ only needs to be large enough. The criteria for GRSSE to be capacity-achieving is significantly easier to satisfy compared to channel coding and lossy source coding.
Even the trivial inner code with $k=0$, $\mathbf{H}=\mathbf{I}$ (where $d=\infty$) can achieve the capacity asymptotically, as (5) gives $C+\log(C+\eta+2)+\eta+3$. In this case, GRSSE reduces to greedy rejection sampling [4]. Nevertheless, using a nontrivial code with a larger $k$ can reduce the running time and the communication cost, since it can reduce the log term in (5). For an AXN channel with bounded noise weight $\mathrm{wt}(\mathbf{Z})\leq w$ (e.g., the Hamming ball channel), we should select a code with $k$ as large as possible while ensuring $d\geq 2w+1$.
Before we present the proof of Theorem 6, we introduce a lemma which may be of interest beyond GRSSE. Unlike greedy rejection sampling (GRS) [4] which has uniquely defined and tractable iterations, the only information we know about the iterations in GRSSE is that $\kappa_{i}$ gives a low rejection probability $\mathbb{E}[\kappa_{i}(\mathrm{e}|\mathbf{S}_{i})]$ (or high acceptance probability) in (2) (the argmin might not even be unique). This flexibility poses a significant challenge to the analysis. The following lemma, which generalizes the bound for GRS in [4], shows that the expected logarithm of the number of iterations can be bounded simply by virtue of having a high acceptance probability for each iteration. It also improves upon [4, 9] by giving a smaller constant. The proof is in Appendix -A.
Lemma 7 (General greedy rejection sampling lemma).
Consider jointly-distributed random variables $Y\in\mathcal{Y}$ and $L\in\mathbb{N}$, a distribution $Q$ over $\mathcal{Y}$ with $P_{Y}\ll Q$, and $\theta>0$. If
[…]
for every $i\in\mathbb{N}$ with $\mathbb{P}(L\geq i)>0$, then
$\mathbb{E}[\log L]\leq D_{\mathrm{KL}}(P_{Y}\|Q)-\log\theta+(1+e^{-1})\log e.$
We remark that for GRS [4], equality holds in (6) for $\theta=1$, so Lemma 7 applies to GRS. Lemma 7 generalizes GRS by allowing the event $L=i$ to be assigned arbitrarily to different values of $y$ (i.e., $\mathbb{P}(L=i|Y=y)$ can be arbitrary), as long as its probability satisfies the lower bound in (6). Lemma 7 also gives a smaller constant $(1+e^{-1})\log e\approx 1.973$ compared to the constant $2\log e\approx 2.885$ in [4, proof of Claim IV.2], and the $\log(2e)\approx 2.443$ in [9]. [^2] We remark that the Poisson functional representation (PFR) satisfies $\mathbb{E}[\log L]\leq D_{\mathrm{KL}}(P_{Y}\|Q)+\log 2$ [27], giving a smaller constant compared to GRS, though PFR is a rather different scheme (e.g., PFR is a noncausal sampling scheme [46] whereas GRS is causal), and does not satisfy (6).
We give a sketch of the proof of Theorem 6. The complete proof is in Appendix -B. We apply Lemma 7 on $(\mathbf{Z},L)$, $\theta=1$,
$Q(\mathbf{z})=\frac{1}{2}\mathbf{1}\Big{\{}\mathrm{wt}(\mathbf{Z})<\frac{d}{2}\Big{\}}q^{k-n}+\frac{1}{2}q^{-n},$
which will yield the desired bounds. [^3] Technically $\sum_{\mathbf{z}}Q(\mathbf{z})\neq 1$. Refer to Appendix -B for the rigorous proof. To check that the acceptance probability at iteration $i$ is at least $F:=\sum_{\mathbf{z}}\min\{p_{\mathbf{Z}}^{(i)}(\mathbf{z}),\,Q(\mathbf{z})\}$ as required by Lemma 7, we need to find one particular $\kappa_{i}(\mathbf{z}|\mathbf{s})$ which gives an acceptance probability at least $F$. Given $\mathbf{s}$, this $\kappa_{i}$ either picks a $\mathbf{z}$ uniformly at random among $\mathbf{z}$’s satisfying $\mathbf{z}\mathbf{H}^{\top}=\mathbf{s}$ with probability $1/2$, or picks the coset leader (the $\mathbf{z}$ giving the smallest $\mathrm{wt}(\mathbf{z})$ with $\mathbf{z}\mathbf{H}^{\top}=\mathbf{s}$) with probability $1/2$. This $\kappa_{i}$ will ensure that a $\mathbf{z}$ with small $\mathrm{wt}(\mathbf{z})$ is selected with enough probability, and hence simulating the part of the noise distribution where the noise weight is small (usually the hardest part to simulate). Since a $\mathbf{z}$ with $\mathrm{wt}(\mathbf{Z})<d/2$ must be a coset leader, we can verify that the flow is indeed at least $F$.
VIII Mixed GRSSE and Experiments
The $\mathbf{H}$-GRSSE in Theorem 6 uses $\mathbf{H}_{i}=\mathbf{H}$ at iteration $i$ which are all the same code. It may be beneficial to allow the inner code $\mathbf{H}_{i}$ to depend on the iteration. Recall that in GRSSE, at iteration $i$, the channel noise distribution to be simulated is $p_{\mathbf{Z}}^{(i)}(\mathbf{z})$, which may change with $i$. Intuitively, in earlier iterations, $p_{\mathbf{Z}}^{(i)}(\mathbf{z})$ is more spread out (i.e., the channel simulated is noisy) so we can use a code $\mathbf{H}_{i}$ with lower rate, or even the trivial code where $k_{i}=0$. In later iterations where $p_{\mathbf{Z}}^{(i)}(\mathbf{z})$ is concentrated around $\mathbf{z}$’s with $\mathrm{wt}(\mathbf{z})\approx 0$ (i.e., we want to send $\mathbf{X}$ almost noiselessly), it is better to use a code with higher rate, or even the complete code where $k_{i}=n$ (i.e., simply send $\mathbf{M}=\mathbf{X}$).
The $\mathbf{H}^{(1)}/\cdots/\mathbf{H}^{(N)}$-GRSSE, or the mixed GRSSE, is the GRSSE where each $\mathbf{H}_{i}$ is selected within a prescribed sequence of codes $\mathbf{H}^{(1)},\ldots,\mathbf{H}^{(N)}$ in this order, i.e., $(\mathbf{H}_{i})_{i}$ is formed by repeating $\mathbf{H}^{(1)}$ a nonnegative integer number of times, then $\mathbf{H}^{(2)}$ a nonnegative integer number of times, etc. The precise number of times can be determined by a computer search that attempts to minimize the expected communication cost. A simple efficient rule that performs well in experiments is to simply select the code $\mathbf{H}^{(j_{i})}$ at iteration $i$ via
$j_{i}:=\max\big{\{}j\in[N]:\,\mathbb{P}_{\mathbf{Z}\sim p_{\mathbf{Z}}^{(i)}}(\mathrm{wt}(\mathbf{Z})\geq\tilde{d}_{j}/2)\leq\epsilon\big{\}},$ (7)
where $\epsilon>0$ is a small fixed parameter, and $\tilde{d}_{j}$ is the distance of the code $\mathbf{H}^{(j)}$ (or an estimate of the distance), assuming that $\tilde{d}_{1}>\cdots>\tilde{d}_{N}$. The reason is that Theorem 6 suggests that a smaller $\mathbb{P}(\mathrm{wt}(\mathbf{Z})\geq d/2)$ would result in a smaller communication cost. This rule shares some similarities with the adaptive greedy rejection sampling scheme [9], which also adaptively changes the proposal distribution at each iteration.
We discuss another reason to consider mixed GRSSE. A limitation of the $\mathbf{H}$-GRSSE is that the precomputation of $\kappa_{i}$ in Algorithm 2 may not terminate, i.e., the number of iterations is unbounded. Even though Algorithm 2 does not terminate, the actual encoding and decoding in Algorithm 1 almost surely terminates since $\mathbb{E}[\log L]<\infty$ by Theorem 6. In practice, we can run Algorithm 2 in a “lazy” manner, where we only compute $\kappa_{i}$ when the encoding algorithm in Algorithm 1 reaches iteration $i$. This way, we can ensure a finite running time for $\mathbf{H}$-GRSSE. Nevertheless, if we require the precomputation in Algorithm 2 to terminate in finite time (e.g., if we want to compute the expected communication cost precisely), we can consider the $\mathbf{H}/\text{complete}$-GRSSE, where $\mathbf{H}^{(1)}=\mathbf{H}$ and $\mathbf{H}^{(2)}$ is the empty matrix (the complete code). When the encoder reachs the iteration where $\mathbf{H}_{i}$ is the complete code, it will simply send $\mathbf{M}=\mathbf{X}$ and terminate the algorithm.
Figure 3 considers the simulation of the memoryless binary symmetric channel with $n=24$ and $\alpha\in[0,1/2]$. The plotted curves are:
  1. 1.
    The Golay-GRSSE algorithm, which is the $\mathbf{H}$-GRSSE with inner code $\mathbf{H}$ being the $[24,12,8]$ Golay code [45]. The expected communication rate (expected encoding length divided by $n$) is $(12+\ell)/n$, where $\ell$ is the expected length of the Huffman encoding of $L$, with $p_{L}$ being computed via Algorithm 2.
  2. 2.
    The mixed-GRSSE algorithm using the rule (7), with $\epsilon=10^{-9}$ and $\mathbf{H}^{(1)},\ldots,\mathbf{H}^{(N)}$ being the best known $[24,k]$ linear block codes [47] with the largest distance for $k=0,1,2,3,5,7,12,14,18,19,23,24$, computed using Magma [48].
  3. 3.
    Poisson functional representation (PFR) [6]. Only the upper bound in [49] is plotted since the blocklength $n=24$ is too large for PFR to be performed efficiently.
  4. 4.
    Greedy rejection sampling (GRS) [4]. Only an upper bound computed via Lemma 7 (which improves upon [4, 9]) is plotted since the blocklength $n=24$ is too large for GRS to be performed efficiently.
  5. 5.
    The capacity of the channel, which is a lower bound on the communication rate.
  6. 6.
    The excess functional information lower bound in [6, Proposition 1], which is a tighter lower bound for finite $n$. (Also see [23, 50] for a generalization.)
Since the Golay code has a distance $8$, according to Theorem 6, Golay-GRSSE is only suitable when $\mathbb{P}(\mathrm{wt}(\mathbf{Z})\leq 3)$ is large, meaning that it requires $\alpha\leq 3/24$. Practically, $\alpha$ has to be lower to ensure $\mathrm{wt}(\mathbf{Z})\leq 3$ most of the time. We can see from Figure 3 that Golay-GRSSE is suitable for $\alpha\leq 0.1$, where its rate is close to the excess functional information lower bound. In comparison, the mixed-GRSSE is close to the lower bound for every $\alpha$, due to the ability to choose a code with a suitable distance at each iteration. The gap between the rate of mixed-GRSSE and the excess functional information lower bound is at most $0.06$ bits.
Figure 4 considers the simulation of the Hamming ball channel with $n=24$ and relative weights $\alpha=w/n$ for $w=0,\ldots,12$, which also serves as lossy source coding under Hamming distortion with almost-sure distortion constraint at distortion level $\alpha$. We plot the same simulation algorithms as Figure 3, and also the best known conventional linear covering codes [17] (note that these codes do not simulate the Hamming ball channels; they merely ensure the same upper bound on the Hamming distortion as the Hamming ball channels). We can see that mixed-GRSSE can outperform conventional linear codes as lossy source codes for distortion between $4/24$ and $10/24$. Mixed-GRSSE also has the advantage of exact channel simulation—the error is exactly uniform over the Hamming ball. This is a similar phenomenon as rejection-sampled universal quantization [11], where channel simulation can also outperform conventional lattice codes in terms of rate-distortion performance.
Figure 3: The communication rate needed by various algorithms to simulate the memoryless binary symmetric channel with $n=24$ and $\alpha\in[0,1/2]$.
Figure 4: The communication rate needed by various algorithms to simulate the Hamming ball channel with $n=24$ and relative weights $\alpha=w/n$ for $w=0,\ldots,12$.
IX GRSSE with Juxtaposition Codes
We now discuss a method to apply GRSSE with a larger blocklength $n$. A simple way to increase the blocklength of a code is to perform juxtaposition, i.e., a codeword in the resultant code is the juxtaposition of any $r$ codewords of the original code. The parity-check matrix of the juxtaposition code is the block-diagonal matrix $\mathbf{I}_{r}\otimes\mathbf{H}$, where where $\mathbf{H}$ is the parity-check matrix of the original code, $\mathbf{I}_{r}$ is the $r\times r$ identity matrix, and $\otimes$ is the Kronecker product. The type set distribution $p_{t_{\mathbf{I}_{r}\otimes\mathbf{H}}}$ (3) of the juxtaposition code $\mathbf{I}_{r}\otimes\mathbf{H}$ can be computed efficiently in terms of the type set distribution $p_{t_{\mathbf{H}}}$ of the original code, by noting that $p_{t_{\mathbf{I}_{r}\otimes\mathbf{H}}}$ is the distribution of $\{r^{-1}\sum_{i=1}^{r}\tau_{i}:\tau_{1}\in t_{1},\ldots,\tau_{r}\in t_{r}\}$ where $t_{i}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}p_{t_{\mathbf{H}}}$, i.e., $p_{t_{\mathbf{I}_{r}\otimes\mathbf{H}}}$ is the convolution (with respect to Minkowski sum) of $r$ copies of $p_{t_{\mathbf{H}}}$.
Figure 5 considers the simulation of the memoryless binary symmetric channel with $n=96$ (except PolarSim) and $\alpha\in[0,1/5]$. The plotted curves are:
  1. 1.
    The 4×4$\times$4 ×Golay-GRSSE with $n=96$, which is the pure GRSSE with inner code $\mathbf{I}_{4}\otimes\mathbf{H}$ being the the juxtaposition of four $[24,12,8]$ Golay codes. [^4] For the sake of efficiency, the number of iterations for both $\mathbf{I}_{4}\otimes\mathbf{H}$-GRSSE and mixed-GRSSE is capped at $20000$, i.e., we switch to the complete code at iteration $20000$.
  2. 2.
    The mixed-GRSSE with $n=96$ using the rule (7), with $\epsilon=0.01$, $\mathbf{H}^{(1)},\ldots,\mathbf{H}^{(N)}$ being the trivial code ($k=0$), the repetition code ($k=1$), and the $4$-fold juxtaposition of the best known $[24,k]$ linear block codes [47] with the largest distance for $k=12,14,18,19,23,24$, computed using Magma [48]. [^5] We choose the better one among this rule and fixing the code to be the $4$-fold juxtaposition of the Golay code (same as the 4×4$\times$4 ×Golay-GRSSE). For the $r$-fold juxtaposition of a code with distance $d_{j}$, we use an estimated distance $\tilde{d}_{j}=rd_{j}$ in rule (7). [^6] Technically, the distance of the $r$-fold juxtaposition is still $d_{j}$. Nevertheless, there are only $r$ codewords with weight $d_{j}$, and their contribution is small in the GRSSE algorithm, which degrades gracefully when a small number of codewords have weights that are too small. Intuitively, the “effective distance” of the $r$-fold juxtaposition is $\tilde{d}_{j}=rd_{j}$, and this estimate performs well in the experiment.
  3. 3.
    PolarSim (channel simulation via polar codes) [12] with $n=4096$.
  4. 4.
    Poisson functional representation (PFR) [6] and greedy rejection sampling (GRS) [4] with $n=96$. Only upper bounds ([49] and Lemma 7) are plotted since the blocklength $n=96$ is exceedingly large for these two schemes to be performed within a year’s time. They are only the hypothetical performance of the impractical schemes, and should not be compared directly to GRSSE which is efficient.
  5. 5.
    The lower bounds given by the capacity of the channel, and the excess functional information lower bound in [6, Proposition 1] with $n=96$.
Figure 5 shows that 4×4$\times$4 ×Golay-GRSSE performs well when $\alpha\in[0.1,0.125]$. We can also see that mixed-GRSSE performs well when $\alpha\leq 0.125$. PolarSim achieves a smaller communication rate, but requires a significantly larger $n$ to have a satisfactory performance. The difference between the rates of mixed-GRSSE ($n=96$) and PolarSim ($n=4096$) is small when $\alpha\leq 0.125$ (the difference is at most $0.058$ bits).
With GRSSE, a scheme with small blocklength but a comparable performance with PolarSim which uses a significantly larger blocklength, we show that a large blocklength is unnecessary to achieve an almost-optimal performance. The applicability of GRSSE under small blocklength allows channel simulation to be performed on shorter sequences with smaller delay.
Figure 5: The communication rate needed by various algorithms to simulate the memoryless binary symmetric channel with $n=96$ (except PolarSim) and $\alpha\in[0,1/5]$. The data for the plot for PolarSim is extracted from [12, Figure 5].
X Acknowledgement
This work was partially supported by two grants from the Research Grants Council of the Hong Kong Special Administrative Region, China [Project No.s: CUHK 24205621 (ECS), CUHK 14209823 (GRF)].